# Prometheus: Infrastructure Security Posture Analysis with AI-generated Attack Graphs

### v1

@@ 8. Evaluation
In this section, we first evaluate the effectiveness of our NER-based model for processing CVE disclosures, followed by an evaluation of our word embedding technique used for attack graph construction. We then present a case study reasoning about the output of Prometheus. Our evaluations aim to answer the following research questions:
@@
We deploy Prometheus on a Google Cloud virtual machine with an Intel Broadwell CPU, 4 GB memory, 110 GB storage, the Ubuntu 20.04 operating system, and an NVIDIA Tesla P100 graphics card.
@@ 8.3. Case Study
Figure 10. Test System for Prometheus
To evaluate Prometheus on capturing the holistic security posture of computer systems, we test it on the infrastructure shown in Figure 10. The test system includes three major components: (1) Jetson Nano device as the server with GPUs, (2) TP-link devices as network switch and router, and (3) Raspberry Pi boards as end devices. Moreover, we also adopt the common communication library, OpenSSL, as the software for securing network traffic.
@@
In the cumulative attack graph, Prometheus has identified CVE-2020-5215, CVE-2020-15206, and CVE-2021-29540 as the top three critical vulnerabilities associated with the installation of two vulnerable Tensorflow versions (2.4.2 and 1.15.2, respectively) on servers equipped with GPUs, as depicted in the spider graph in Figure 8. These vulnerabilities are notably part of the computed vertex cover, indicating their presence in the set of vulnerabilities encompassing all attack paths. Additionally, Prometheus highlighted the top three attack paths based on their elevated risk, exploitability, and impact, as shown in the “top path risk vectors” section of Figure 8. It is worth noting that in every generated attack graph, at least one node corresponds to the computed vertex cover, identifying the minimum set of vulnerabilities that must be addressed. Addressing these vulnerabilities through patching may render all attack graphs impractical. Ultimately, Prometheus has identified the five shortest attack paths within the cumulative attack graph, each having one CVE node with an exploitability score of 10, denoting the highest possible score. These CVE nodes are specifically CVE-2010-3173 (Mozilla Firefox on end devices), CVE-2011-0392 (Cisco TelePresence Recording Server on the server), CVE-2012-6531 (Zend Framework on end devices), CVE-2012-0884 (OpenSSL in all devices and the server), and CVE-2015-0763 (Cisco Unified MeetingPlace on end devices). Notably, analogous observations hold true for the network, system & hardware, machine learning, and cryptography layers.
@@ 9. Related Work
Risk Scoring. Li et al. (li2006cluster, ) have proposed a cost/benefit analysis of the attack graphs. Administrators have to manually provide cost and benefit values for every node in the attack graph. A cost may relate to the adversary in terms of the resources that they need to possess, whereas a benefit is for the organization. Lu et al. (lu2009ranking, ) use the graph neural networks (GNNs) model to rank attack graphs. The authors use Google’s PageRank web search algorithm (bianchini2005inside, ) to infer each step’s importance in the attack graph. Prometheus, however, extends the CVSS scoring system, which therefore aligns with the assessment and expertise of the CVE assigner. Idika et al. (idika2010extending, ) have proposed attack graph-based security metrics to analyze attack graphs. However, their analysis is focused on the shortest paths and the number of paths an attacker can take to violate a security policy and not on the severity of the attack paths.
@@ 10. Conclusion and Future Work
The unceasing growth of cybercrime and network complexity require innovative cybersecurity solutions. Prometheus, the security analyzer proposed in this paper, is a promising step toward meeting this demand. Through the use of machine learning, natural language processing, and systematic vulnerability analysis, Prometheus offers a comprehensive security solution for general computing infrastructures in that it is capable of extracting semantic meanings from disclosed vulnerabilities, performing layered classification, and creating exhaustive directed attack graphs for holistic security analytics.
The distinctive multi-layered approach to security posture analysis provided by Prometheus holds the potential for a more nuanced understanding of a network’s security landscape, allowing vulnerabilities of different natures and their interrelations to be effectively addressed. Furthermore, the proposed scoring methods provide a more tailored and relevant security assessment for the infrastructure under analysis, thereby enhancing the practical value of the analytics.
While the application of Prometheus shows promise, there is significant scope for future work. First, improving the machine learning models used in Prometheus will be crucial. As these models evolve, the accuracy and effectiveness of semantic extraction of vulnerabilities can be enhanced, leading to a more reliable and detailed vulnerability analysis. Second, enhancing the granularity of the layer classification process is an interesting research direction. While Prometheus classifies vulnerabilities into layers, refining this classification to include sub-layers or specialized categories might provide more targeted insights for security measures. Third, developing advanced scoring methods can be a pivotal area of future exploration. As Prometheus traverses the generated graph, devising innovative and more granular metrics to gauge the severity of vulnerabilities could make risk assessment more precise.
Furthermore, Prometheus could be extended to incorporate automated patching and response capabilities to enhance its practical utility. This would allow for immediate remediation of identified vulnerabilities, enhancing the network’s overall security. Integrating Prometheus with real-time monitoring systems could also provide dynamic, up-to-date security posture analytics, thereby enabling faster response times to potential threats. Last, real-world testing of Prometheus across various industries, like healthcare, energy, or finance, would provide invaluable feedback for fine-tuning the system. These tests could also uncover specific industry-related vulnerabilities, which could be incorporated into the model for a more sector-specific approach.
Overall, Prometheus offers significant strides in cybersecurity. Still, its evolution is a continuous journey that will shape the contours of a safer and more secure digital landscape.
@@ Appendix A Keywords and CWE Lists for layer matching
In this section, we list a subset of keywords used by Prometheus to classify the vulnerabilities to one of our proposed layers; the Network Security layer. Table 6 lists the network-related keywords, whereas Table 7 lists a subset of widely adopted communication protocols in network infrastructures.

### v2

@@ 8. Evaluation
In this section, we first evaluate the effectiveness of our NER-based model for processing CVE disclosures, followed by an evaluation of our word embedding technique used for attack graph construction. We then present a case study reasoning about the output of Graphene. Our evaluations aim to answer the following research questions:
@@
We deploy Graphene on a Google Cloud virtual machine with an Intel Broadwell CPU, 4 GB memory, 110 GB storage, the Ubuntu 20.04 operating system, and an NVIDIA Tesla P100 graphics card.
@@ 8.3. Case Study
Figure 10. Test System for Graphene
To evaluate Graphene on capturing the holistic security posture of computer systems, we test it on the infrastructure shown in Figure 10. The test system includes three major components: (1) Jetson Nano device as the server with GPUs, (2) TP-link devices as network switch and router, and (3) Raspberry Pi boards as end devices. Moreover, we also adopt the common communication library, OpenSSL, as the software for securing network traffic.
@@
In the cumulative attack graph, Graphene has identified CVE-2020-5215, CVE-2020-15206, and CVE-2021-29540 as the top three critical vulnerabilities associated with the installation of two vulnerable Tensorflow versions (2.4.2 and 1.15.2, respectively) on servers equipped with GPUs, as depicted in the spider graph in Figure 8. These vulnerabilities are notably part of the computed vertex cover, indicating their presence in the set of vulnerabilities encompassing all attack paths. Additionally, Graphene highlighted the top three attack paths based on their elevated risk, exploitability, and impact, as shown in the “top path risk vectors” section of Figure 8. It is worth noting that in every generated attack graph, at least one node corresponds to the computed vertex cover, identifying the minimum set of vulnerabilities that must be addressed. Addressing these vulnerabilities through patching may render all attack graphs impractical. Ultimately, Graphene has identified the five shortest attack paths within the cumulative attack graph, each having one CVE node with an exploitability score of 10, denoting the highest possible score. These CVE nodes are specifically CVE-2010-3173 (Mozilla Firefox on end devices), CVE-2011-0392 (Cisco TelePresence Recording Server on the server), CVE-2012-6531 (Zend Framework on end devices), CVE-2012-0884 (OpenSSL in all devices and the server), and CVE-2015-0763 (Cisco Unified MeetingPlace on end devices). Notably, analogous observations hold true for the network, system & hardware, machine learning, and cryptography layers.
@@ 9. Related Work
Risk Scoring. Li et al. (li2006cluster, ) have proposed a cost/benefit analysis of the attack graphs. Administrators have to manually provide cost and benefit values for every node in the attack graph. A cost may relate to the adversary in terms of the resources that they need to possess, whereas a benefit is for the organization. Lu et al. (lu2009ranking, ) use the graph neural networks (GNNs) model to rank attack graphs. The authors use Google’s PageRank web search algorithm (bianchini2005inside, ) to infer each step’s importance in the attack graph. Graphene, however, extends the CVSS scoring system, which therefore aligns with the assessment and expertise of the CVE assigner. Idika et al. (idika2010extending, ) have proposed attack graph-based security metrics to analyze attack graphs. However, their analysis is focused on the shortest paths and the number of paths an attacker can take to violate a security policy and not on the severity of the attack paths.
@@ 10. Conclusion and Future Work
The unceasing growth of cybercrime and network complexity require innovative cybersecurity solutions. Graphene, the security analyzer proposed in this paper, is a promising step toward meeting this demand. Through the use of machine learning, natural language processing, and systematic vulnerability analysis, Graphene offers a comprehensive security solution for general computing infrastructures in that it is capable of extracting semantic meanings from disclosed vulnerabilities, performing layered classification, and creating exhaustive directed attack graphs for holistic security analytics.
The distinctive multi-layered approach to security posture analysis provided by Graphene holds the potential for a more nuanced understanding of a network’s security landscape, allowing vulnerabilities of different natures and their interrelations to be effectively addressed. Furthermore, the proposed scoring methods provide a more tailored and relevant security assessment for the infrastructure under analysis, thereby enhancing the practical value of the analytics.
While the application of Graphene shows promise, there is significant scope for future work. First, improving the machine learning models used in Graphene will be crucial. As these models evolve, the accuracy and effectiveness of semantic extraction of vulnerabilities can be enhanced, leading to a more reliable and detailed vulnerability analysis. Second, enhancing the granularity of the layer classification process is an interesting research direction. While Graphene classifies vulnerabilities into layers, refining this classification to include sub-layers or specialized categories might provide more targeted insights for security measures. Third, developing advanced scoring methods can be a pivotal area of future exploration. As Graphene traverses the generated graph, devising innovative and more granular metrics to gauge the severity of vulnerabilities could make risk assessment more precise.
Furthermore, Graphene could be extended to incorporate automated patching and response capabilities to enhance its practical utility. This would allow for immediate remediation of identified vulnerabilities, enhancing the network’s overall security. Integrating Graphene with real-time monitoring systems could also provide dynamic, up-to-date security posture analytics, thereby enabling faster response times to potential threats. Last, real-world testing of Graphene across various industries, like healthcare, energy, or finance, would provide invaluable feedback for fine-tuning the system. These tests could also uncover specific industry-related vulnerabilities, which could be incorporated into the model for a more sector-specific approach.
Overall, Graphene offers significant strides in cybersecurity. Still, its evolution is a continuous journey that will shape the contours of a safer and more secure digital landscape.
@@ Appendix A Keywords and CWE Lists for layer matching
In this section, we list a subset of keywords used by Graphene to classify the vulnerabilities to one of our proposed layers; the Network Security layer. Table 6 lists the network-related keywords, whereas Table 7 lists a subset of widely adopted communication protocols in network infrastructures.